# Confidence-based Estimators for Predictive Performance in Model Monitoring

### v1

@@ 7 Conclusion
In this study, we revisited AC, a commonly used baseline method in unsupervised accuracy estimation. We analyzed its theoretical properties under the assumption of perfect calibration and showed it to be an unbiased and consistent estimator of the predictive performance of a deployed ML model. We showed experimentally that by leveraging the properties of the Poisson binomial distribution, it is possible to find valid confidence intervals for the estimates provided by AC, which is a property no other confidence-based estimator currently has to the best of our knowledge. We compared the quality of the estimates provided by AC against more advanced confidence-based estimators and experimentally showed that despite its simplicity, it is competitive against other estimators, as long as the calibration error remains small enough.
@@ Appendix A. Data description for the covariate shift experiment
Data used in the first scenario
@@
Data used in the second scenario
@@
Contrary to the first scenario, this way of sampling induced a label shift, with an increasing fraction of positive labels as shown in Table 8. However, we did not consider this to be a source of error given that the model performance deterioration rate was very similar to what was observed in the case of a linear decision boundary.
Table 8: The fraction of positive labels in the second scenario.

### v2

@@ 7 Conclusion
In this study, we revisited AC, a commonly used baseline method in unsupervised accuracy estimation. We analyzed its theoretical properties under the assumption of perfect calibration and showed it to be an unbiased and consistent estimator of the predictive performance of a deployed ML model. We showed experimentally that by leveraging the properties of the Poisson binomial distribution, it is possible to find valid confidence intervals for the estimates provided by AC, which is a property no other confidence-based estimator currently has to the best of our knowledge. We compared the quality of the estimates provided by AC against more advanced confidence-based estimators and experimentally showed that despite its simplicity, AC is competitive against these estimators, as long as the calibration error remains small enough. In addition to potential calibration error, encountering concept shift or data outside the support of training data might hamper the performance of AC when monitoring a model in production. Extending the scope of confidence-based performance estimation to include other metrics in addition to accuracy and comparing the estimators in real-life scenarios are important challenges, which we leave for future research.
Acknowledgments
This work was partly funded by local authorities (“Business Finland”) under grant agreement 20219 IML4E of the ITEA4 programme.
@@ Appendix A. Data Description for the Covariate Shift Experiment
A1. Data Used in the First Scenario
@@
A2. Data Used in The Second Scenario
@@
Contrary to the first scenario, this way of sampling induced a label shift, with an increasing fraction of positive labels as shown in Table 8.
@@
Table 8: The fraction of positive labels in the second scenario.
However, we did not consider this to be a source of error given that the model performance deterioration rate was very similar to what was observed in the case of a linear decision boundary.